1. Four identical lamps are connected in parallel to a 6 volt battery. What is the voltage drop across each lamp?
  2. What happens to the total resistance in a network when more resistors are added to a series circuit ?
  3. Four resistors of resistances 3Ω , 4Ω , 5Ω and 6Ω are connected in parallel to a 12 volts source, what is the current through the 5Ω resistor.
1
Expert's answer
2021-06-10T08:46:23-0400

Gives four identical lamp resistance "R=R_1=R_2=R_3=R_3=R_4"

All resistance is parallel combination

"R_{net}=\frac{R}{4}"

Voltage drop across each resistance are equal

"V=" 6V

"=V=V_1=V_2=V_3=V_4=6V"

(2)

Four resistance is series combination

"R_{net}=R_1+R_2+R_3+R_4"

Rnet =3+4+5+6=14 "\Omega"

(3)

All four resistance are parallel combination

"\frac{1}{R}=\frac{20+15+12+10}{60}=\frac{57}{60}"

"R=\frac{60}{57}=1.05\Omega"

V=12V

"I=\frac{12}{1.05}=11.4A"

R=5ohm ResistanceCurrent through the find out

Ratio of current



"I_1:I_2:I_3:I_4=\frac{1}{R_1}:\frac{1}{R_2}:\frac{1}{R_3}:\frac{1}{R_4}:"

Put Value of all resistance

Ratio of currents

"I_1:I_2:I_3:I_4=20:15:12:10"


"I_3=\frac{12}{57}\times11.4" =2.4A